OpenStudy (anonymous):

please help!!! f(1) = 2 and f(2) = 5, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______

OpenStudy (anonymous):

Will give medal and fan!!

OpenStudy (anonymous):

this is what i got F(x) = f(x f(x f(x))) F‘(x) = f‘(x f(x f(x))) * [f(x f(x)) + x {f‘(x f(x)) * [f(x) + x f‘(x)]}] F '(1) = f‘(f(f(1))) * [f(f(1)) + {f‘(f(1)) * [f(1) + f‘(1)]}] F '(1) = f‘(f(2)) * [f(2) + {f‘(2) * [f(1) + 1]}] F '(1) = f‘(3) * [3 + {2 * [2 + 1]}] F '(1) = 3 * [3 + {2 * [2 + 1]}] F '(1) = 27 -

OpenStudy (anonymous):

sorry i dont like when that happens

OpenStudy (academicgurusinc):

F(5) = f(1) + f(2) + f(5-1)= 2 + 5 + f(4) =7 + f(4) now find f(4): f(4) = 2 + 5 + f(3)= 7 + f(3) now find f(3): f(3) = 2 + 5 + f(2) = 7 + 5= 12 now solve f(4): f(4) = 7+ 12 =19 now solve f(5): f(5) = 7 + 19 = 26 We hope this helps. Regards, Academic GurusInc. Twitter (@Academic_Gurus) Facebook (AcademicGurusInc)

OpenStudy (anonymous):

thx thats what i was trying to say

OpenStudy (academicgurusinc):

Hello thegeneral, Your question was the featured question on our Tutoring Tuesday session. Link: https://www.youtube.com/watch?v=xi6NAS13Xrg Give us a thumbs up or subscribe if you liked the video or submit suggestions to questions@academicgurusinc.com We'd love to hear from you. Best Regards, AcademicGurusInc